Convert 912 Torr to atmospheres Convert 28.5 atmospheres to torr. Torr atm Convert 912 mmHg to atmospheres. Convert 760 Torr to millimeters of mercury atm mmHg Previous Ove Up & View Solution Check Answer NextEx Hint MacBook Pro F8 F7 F6 20 ao pa FS 3
Answer #1

1) 1 torr = 0.00132 atm

912 torr = 0.00132× 912 atm

= 1.203 atm

2) 0.00132 atm = 1 torr

28.5 atm = 28.5/0.00132 torr

= 21590.9 torr

3) 760 mmHg = 1atm

912 mmHg = 912/760 atm

= 1.2 atm

4)1 torr = 0.00132 atm

760 torr= 0.00132 × 760 atm

=1.0032 atm

1 atm = 760mmHg

1.0032 atm = 760 × 1.0032 mmHg

= 762.43 mmHg